OpenStudy (anonymous):

A cubic meter is 1000000 times as large as a cubic decimeter? Is that correct....

OpenStudy (anonymous):

1 m^3 = 10^3 dm^3 = 1000 dm^3

OpenStudy (anonymous):

Huh

OpenStudy (anonymous):

@ganeshie8